OpenStudy (cutiecomittee123):

The angle of elevation to the top of a skyscraper is measured to be 2 degrees from a point on the ground 1 mile from the building. How tall is the skyscraper?

OpenStudy (jacob902):

Well, I guess you start with a triangle. Let O be the point 1 mile from the building. Let B be the bottom of the building Let H be the top of the building. Angle HOB is given (2 degrees) Side OB is given as 1 mile (5280 feet) So side BH = the height of the building tan(HOB) = BH / OB BH = OB * tan(HOB)

OpenStudy (jacob902):

| | h | |________________________2°__ e = 1 mile tan 2° = h/e h = e * tan 2° h = 0.0349 miles = 61,46 yard

OpenStudy (jacob902):

0.0349207695 mi

OpenStudy (cutiecomittee123):

you lost me after tan(2)=x/1

OpenStudy (anonymous):

 h=184.4 ft.

OpenStudy (cutiecomittee123):

got it, thanks:)